A $150,000 mortgage is amortized over 25 years. If interest on the mortgage is 3.5 percent compounded semi-annually, calculate the size of monthly payments made at the end of each month.

A. $784.91

B. $748.91

C. $734.91

D. $743.91
Answer #1

here n= 2, Number of months p = 6
Interest rate per month =(1+i/n)n/p-1 = (1+3.5%/2)1/12-1 = 0.2896%

Monthly payments = PV/(1-(1+r)-n)/r = 150,000/((1-(1+0.2896%)-12*25)/0.2896% = 748.91

Option b is correct option

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well